OpenStudy (anonymous):

the sideof a square is(x-5cm.what is area

OpenStudy (anonymous):

area would simply be (x-5)(x-5) = [(x^2)-(2)(x)(5)+(5^2)] = X^2-10x+25 This is because the sides of a square are equal. Therefore when you want to find the area of a square, always just square the length/breath.

OpenStudy (marissalovescats):

Also, welcome to OpenStudy! @hans1234